# VD3D: Taming Large Video Diffusion Transformers for 3D Camera Control

### v2

@@ 4.2 Ablations
Plucker embedding. We motivate our Plucker embedding conditioning mechanism by training a variant using the raw camera matrices. Concretely, we flatten and concatenate extrinsics and intrinsics matrices in the channel dimension and repeat the values in the spatial patch dimensions. We observe that Plucker embeddings provide an essential spatial conditioning mechanism, as shown in Tab. 2.
ControlNet conditioning. Our ControlNet inspired conditioning mechanism ensures fast and precise learning of the conditioning signal distribution. Instead of using a ControlNet block, we simply add zero-initialized Plucker embedding features to the patches and observe close to no camera control. We observe training cross-attention layers in the ControlNet block is key to learning camera control while preserving the original model weights. This is confirmed by our camera evaluation in Tab. 2.
@@ 4.3 Applications
Image-to-video generation. We synthesize camera-controlled videos based on different camera poses as shown in Fig. 1. For this task we use a version of the SnapVideo model that we fine-tune on video sequences where a random subset of the input frames are masked. At inference time, we can provide image guidance for any of the generated frames, providing an additional dimension of controllability when paired with our proposed method for camera control. To demonstrate image-to-video generation in Fig. 1, we condition the model using camera poses along with image guidance from the first frame of a generated video sequence. Note that our method provides no control over motion within the scene itself; hence, scene motion can differ depending on the random seed or the provided camera poses.
Image-to-multiview generation. We also explore multi-view generation for static scenes as shown in Fig. 5. Given a real input image of a complex scene unseen during training, our camera-conditioned model generates view-consistent renderings of that scene from arbitrary viewpoints. These multi-view renderings could be directly relevant to downstream 3D reconstructions pipelines, e.g., based on NeRF [74] or 3D Gaussian Splatting [49]. We show the potential of camera-conditioned image-to-multiview generation for complex 3D scene generation, but we leave more extensive exploration of this topic for future work.
@@ 5 Conclusion
Broader impacts.
@@ Appendix A Related Work
Limited to single category trainings, early work on 3D generation extends GANs into 3D using a neural renderer as an inductive bias [18, 10, 76, 89, 3], Towards more diverse and flexible generation, CLIP-based supervision [84] enabled text-based generation and editing of 3D assets [14, 42, 88, 43, 25, 113]. With recent advances in diffusion models, Score Distillation Sampling (SDS) replaces CLIP supervision with diffusion model supervision  [80, 118, 61, 15, 60, 114, 58, 30, 138, 68, 144, 48, 53, 96] for higher quality generation. In order to improve the 3D structure of scenes, another line of work generates multiple views of a scene [62, 67, 91, 20, 66, 50, 111, 35]. Alternatively, other methods use iterative inpainting for 3D scene generation [36, 92]. Recent methods lift input images into 3D  [11, 102, 26, 69, 140, 103, 83, 70, 112, 99] using NeRF [74], 3D Gaussian Splatting [49], or Meshes in combination with diffusion models. Other recent methods [37, 56, 131, 130, 147, 28, 44, 126, 101, 104, 81, 100, 98] tackle fast feed-forward 3D generation, directly predicting a 3D generation from input images or text. Different from our approach, these methods can only synthesize static scenes.
@@ B.1 Multi-view generation
We evaluate our model for image-to-multiview generation. Due to ground truth correspondences for the RealEstate10K [156] dataset, we can condition our model on a given camera trajectory and assess image based metrics, i.e., PSNR, SSIM, and LPIPS. We provide results for the low-resolution base model and the upsampled high-resolution results in Tab. 3.
@@ B.2 Quality metrics
Moreover, we evaluate all models on established image and video generation metrics, namely, FID [32], FVD [106], and CLIPSIM [122]. We provide results for RealEstate10K [156] and MSR-VTT [129] in Tab. 4.

### v3

@@ 4.2 Ablations
Plücker embedding. We motivate our Plücker embedding conditioning mechanism by training a variant using the raw camera matrices. Concretely, we flatten and concatenate extrinsics and intrinsics matrices in the channel dimension and repeat the values in the spatial patch dimensions. We observe that Plücker embeddings provide an essential spatial conditioning mechanism, as shown in Tab. 2.
ControlNet conditioning. Our ControlNet-inspired conditioning mechanism ensures fast and precise learning of the conditioning signal distribution. Instead of using a ControlNet block, we simply add zero-initialized Plücker embedding features to the patches and observe close to no camera control. We observe training cross-attention layers in the ControlNet block is key to learning camera control while preserving the original model weights. This is confirmed by our camera evaluation in Tab. 2.
@@
Add to context vector. We train a model where we use camera matrices, linearly encode them, and add them to the context vector as a simple conditioning mechanism for transformer-based models. The rendered videos show little to no correlation with the input camera matrices, resulting in poor camera pose accuracy, as shown in Tab. 2. This highlights the importance of carefully incorporating cameras into the spatio-temporal patches with our Plücker embedding block.
Plücker in context vector. We train a model where we integrate Plücker embeddings into the context vector instead of the patches. We similarly compute Plücker embeddings, but instead of patchifying them, we flatten and linearly map the embeddings into a vector that has the same shape as the context vector. We fuse the Plücker features and context vector using the same Plücker conditioning mechanism used for the patches. We observe higher camera pose errors in Tab. 2, highlighting that it is crucial to incorporate the spatio-temporal Plücker embeddings into the spatio-temporal patches instead of using the context vector.
@@ 4.3 Applications
Image-to-video generation. We synthesize camera-controlled videos based on different camera poses as shown in Fig. 1. For this task we use a version of our pre-trained text-to-video model that we fine-tune on video sequences where a random subset of the input frames are masked. At inference time, we can provide image guidance for any of the generated frames, providing an additional dimension of controllability when paired with our proposed method for camera control. To demonstrate image-to-video generation in Fig. 1, we condition the model using camera poses along with image guidance from the first frame of a generated video sequence. Note that our method provides no control over motion within the scene itself; hence, scene motion can differ depending on the random seed or the provided camera poses.
Image-to-multiview generation. We also explore multi-view generation for static scenes as shown in Fig. 5. Given a real input image of a complex scene unseen during training, our camera-conditioned model generates view-consistent renderings of that scene from arbitrary viewpoints. These multi-view renderings could be directly relevant to downstream 3D reconstructions pipelines, e.g., based on NeRF (Mildenhall et al., 2020) or 3D Gaussian Splatting (Kerbl et al., 2023). We show the potential of camera-conditioned image-to-multiview generation for complex 3D scene generation, but we leave more extensive exploration of this topic for future work.
@@ 5 Conclusion
6 Ethics Statement
Broader Impact.
@@ 6 Ethics Statement
Data.
To develop the camera control methods proposed in this paper, we used RealEstate10K (Zhou et al., 2018). RealEstate10K is released and open-sourced by Google LLC under a Creative Commons Attribution 4.0 International License, and sourced from content using a CC-BY license. The dataset can be found under the following URL: https://google.github.io/realestate10k. The RealEstate10K dataset was published as part of a research paper by Zhou et al. (2018): “Stereo Magnification: Learning View Synthesis using Multiplane Images" at SIGGRAPH 2018. The RealEstate10K dataset is a commonly used dataset for 3D reconstruction (Charatan et al., 2024), 3D generation (Gao et al., 2024c), and camera-controlled video generation (Wang et al., 2023e). Similarly, closely related baselines such as MotionCtrl (Wang et al., 2023e) and CameraCtrl (He et al., 2024a) use the identical dataset to train and evaluate their camera-controlled video diffusion models. Hence, we follow an established training and evaluation pipeline without any modifications to the data.
7 Reproducibility Statement
We have structured our paper to ensure comprehensive reproducibility of our camera control method. Section 3 provides a detailed description of our method, including theoretical foundations and core algorithmic components. Section 4 presents thorough experimental details, covering evaluation protocols, metrics, and comparisons with baselines. All implementation specifics, including training procedures, architectural choices, and hyperparameters, are documented in Appendix C. To further facilitate reproducibility, we include the source code of our camera-controlled FIT block as supplementary material. The provided implementation contains the core components necessary to replicate our approach, accompanied by documentation and usage examples. We welcome requests for additional technical details or clarifications to support the research community in reproducing and building upon our work.
8 Acknowledgements
DBL acknowledges support from NSERC under the RGPIN program, the Canada Foundation for Innovation, and the Ontario Research Fund.
@@ Appendix A Related Work
Due to space constraints, we summarize related 3D and an extended list of 4D works in the appendix.
@@
Limited to single category trainings, early work on 3D generation extends GANs into 3D using a neural renderer as an inductive bias (DeVries et al., 2021; Chan et al., 2022; Or-El et al., 2022; Schwarz et al., 2022; Bahmani et al., 2023b), Towards more diverse and flexible generation, CLIP-based supervision (Radford et al., 2021) enabled text-based generation and editing of 3D assets (Chen et al., 2018; Jain et al., 2022; Sanghi et al., 2022; Jetchev, 2021; Gao et al., 2023; Wang et al., 2022). With recent advances in diffusion models, Score Distillation Sampling (SDS) replaces CLIP supervision with diffusion model supervision  (Poole et al., 2023; Wang et al., 2023d; Lin et al., 2023a; Chen et al., 2023c; Liang et al., 2023; Wang et al., 2023a; Li et al., 2024d; He et al., 2024b; Ye et al., 2024; Liu et al., 2024c; Yu et al., 2023b; Katzir et al., 2024; Lee et al., 2024a; Sun et al., 2024a) for higher quality generation. In order to improve the 3D structure of scenes, another line of work generates multiple views of a scene (Lin et al., 2023b; Liu et al., 2023; Shi et al., 2024; Feng et al., 2024a; Liu et al., 2024b; Kim et al., 2023; Voleti et al., 2024b; Höllein et al., 2024). Alternatively, other methods use iterative inpainting for 3D scene generation (Höllein et al., 2023; Shriram et al., 2024). Recent methods lift input images into 3D  (Chan et al., 2023; Tang et al., 2023; Gu et al., 2023; Liu et al., 2024d; Yoo et al., 2023; Tewari et al., 2023; Qian et al., 2024c; Long et al., 2024; Wan et al., 2024; Szymanowicz et al., 2023) using NeRF (Mildenhall et al., 2020), 3D Gaussian Splatting (Kerbl et al., 2023), or Meshes in combination with diffusion models. Other recent methods (Hong et al., 2024; Li et al., 2024b; Xu et al., 2024d; c; Zhang et al., 2024a; Han et al., 2024; Jiang & Wang, 2024; Xie et al., 2024; Tang et al., 2024; Tochilkin et al., 2024; Qian et al., 2024a; Szymanowicz et al., 2024b; a) tackle fast feed-forward 3D generation, directly predicting a 3D generation from input images or text. Different from our approach, these methods can only synthesize static scenes.
4D generation. Previous methods also tackle the problem of 4D generation, i.e., generating videos of dynamic 3D scenes from controllable viewpoints, usually from an input text prompt or image. Since the initial work on this topic using large-scale generative models (Singer et al., 2023b), significant improvements in the visual quality and motion quality of generated scenes have been achieved (Ren et al., 2023; Ling et al., 2024a; Bahmani et al., 2024b; Zheng et al., 2024a; Gao et al., 2024b; Yang et al., 2024a; Jiang et al., 2024; Bahmani et al., 2024a; Zhang et al., 2024d; Xu et al., 2024a; Miao et al., 2024; Li et al., 2024a). While these methods generate scenes based on text conditioning, other approaches convert an input image or video to a dynamic 3D scene (Ren et al., 2023; Zhao et al., 2023; Yin et al., 2023; Pan et al., 2024; Zheng et al., 2024a; Ling et al., 2024a; Gao et al., 2024b; Zeng et al., 2024; Chu et al., 2024; Wu et al., 2024; Yang et al., 2024c; Wang et al., 2024; Feng et al., 2024b; Sun et al., 2024b; Zhang et al., 2024b; Yu et al., 2024; Ren et al., 2024; Lee et al., 2024b; Li et al., 2024c; Van Hoorick et al., 2024; Uzolas et al., 2024; Huang et al., 2024; Chai et al., 2024; Liang et al., 2024; Zhang et al., 2024c). Another line of work (Bahmani et al., 2023a; Xu et al., 2023) extends 3D GANs into 4D by training on 2D videos, however the quality is limited and models are trained on single category datasets. Still, all of these methods are focused on object-centric generation, typically based on 3D volumetric representations. As such, they typically do not incorporate background elements, and overall, they do not approach the level of photorealism demonstrated by the state-of-the-art video generation models used in our technique.
@@ B.1 Multi-view generation
We evaluate our model for image-to-multiview generation. Due to ground truth correspondences for the RealEstate10K (Zhou et al., 2018) dataset, we can condition our model on a given camera trajectory and assess image based metrics, i.e., PSNR, SSIM, and LPIPS. We provide results for the low-resolution base model and the upsampled high-resolution results in Tab. 3.
@@ B.2 Quality metrics
Moreover, we evaluate all models on established image and video generation metrics, namely, FID (Heusel et al., 2017), FVD (Unterthiner et al., 2018), and CLIPSIM (Wu et al., 2021). We provide results for RealEstate10K (Zhou et al., 2018) and MSR-VTT (Xu et al., 2016) in Tab. 4.
It is important to highlight that FID and FVD are not direct measurements of visual quality—these metrics measure similarities between dataset distributions. So while fine-tuning on a dataset will degrade these scores relative to the original model, this degradation is expected because the model is fitting a different data distribution than the one used in the original training run. The same degradation in FID after fine-tuning is observed in the original ControlNet paper (Zhang et al., 2023), where the authors report an increase in FID when comparing the StableDiffusion base model (6.09 FID) to their fine-tuned ControlNet model (15.27 FID; see their Table 3). We believe that joint training with 2D video data may help to alleviate this degradation, and we will explore joint training strategies as part of future work.
B.3 Baseline variants
We provide further variants of MotionCtrl and CameraCtrl adopted to the SnapVideo base model. These variants performed worse than the ones presented in the main paper, hence we mainly show them here for completeness to show the different implementations we explored.
On top of that, we train a variant of MotionCtrl where instead of integrating the camera matrices into the context vector, we integrate them into the output of the cross-attention layer between the latent tokens and patches (as in our approach). We observe degraded camera control compared to our approach. Moreover, we train a variant of CameraCtrl where we incorporate the Plucker embeddings of the CameraCtrl camera encoder into the patches instead of the latents. We observe worse camera control compared to our model. Furthermore, we train a variant of MotionCtrl where we freeze all base model layers similar to CameraCtrl and our approach. We use the same batch size, same number of iterations, and same parameter size as our proposed method for fair comparisons. Note that MotionCtrl unfreezes the attention layer after injecting camera features into the base model. This experiment highlights that we outperform MotionCtrl independent of training or freezing these attention layers. We show results for camera pose accuracy, multi-view generation, and video quality in Tab. 5, Tab. 6, and Tab. 7 respectively.
B.4 Generalization of camera trajectories
We conduct additional experiments with horizontal and vertical panning, where the camera trajectory is defined by rotation-only camera matrices without any translation. Specifically, we manually construct each trajectory by randomly selecting either the x-, y-, or z-axis and randomly sampling the angular extent of the trajectory (from 0 to 120 degrees). We set the camera translation to the zero vector. We show results in Tab. 8 and observe higher accuracy for our methods. Furthermore, we include results for non-random, user-defined camera trajectories that involve camera movements with significant directional changes including both rotations and translations in Tab. 9. These include following trajectories: rotation around clockwise; rotation around anticlockwise; rotation clockwise without translation; rotation anticlockwise without translation; zoom out, then up translation; translation right, then rotation anticlockwise; translation left, then rotation clockwise; translation left; translation right; translation up; translation down. We observe that our method generalizes to input camera trajectories with variable rotations and translations.
B.5 Experiments with vanilla DiT
Instead of building upon the FIT (Chen & Li, 2023) architecture, we implemented our VD3D method on top of a pre-trained text-to-video DiT model (Peebles & Xie, 2023) in the latent space of the CogVideoX (Yang et al., 2024d) autoencoder. We include these results in Tab. 10 and Tab. 11. Instead of building on top of read attention in FIT, we incorporate the ControlNet conditioning on top of the vanilla attention mechanism of the actual tokens. We observe that the vanilla DiT version further improves quality and camera accuracy on out-of-distribution prompts (MSR-VTT). We believe that our proposed method of spatio-temporal Plucker tokens and aligning them with video patch tokens through a ControlNet-type of conditioning mechanism is agnostic to the transformer architecture and will serve as a starting point for follow-up works.
Base video DiT architecture details.
The video DiT architecture follows the design of other contemporary video DiT models (e.g., Sora (Brooks et al., 2024), MovieGen (Polyak et al., 2024), OpenSora (Zheng et al., 2024b), LuminaT2X (Gao et al., 2024a), and CogVideoX (Yang et al., 2024d)). As the backbone, it incorporates a transformer-based architecture with 32 DiT blocks. Each DiT block includes a cross-attention layer for processing text embeddings (produced by the T5-11B model), a self-attention layer, and a fully connected network with a ×4 dimensionality expansion. Attention layers consist of 32 heads with RMSNorm for query and key normalization. Positional information is encoded using 3D RoPE attention, where the temporal, vertical, and horizontal axes are allocated fixed dimensionality within each attention head (using a 2:1:1 ratio). LayerNorm is applied to normalize activations within each DiT block. A pre-trained CogVideoX autoencoder is utilized for video dimensionality reduction, employing causal 3D convolution with a 4×8×8 compression rate and 16 channels per latent token. The model features a hidden dimensionality of 4,096 and comprises 11.5B parameters. It leverages block modulations to condition the video backbone on rectified flow timestep information, SiLU activations, and 2×2 ViT-like patchification of input latents to reduce sequence size.
Base video DiT training details.
The base DiT model is optimized using AdamW, with a learning rate of 0.0001 and weight decay of 0.01. It is trained for 750,000 iterations with a cosine learning rate scheduler in bfloat16. Image animation support is incorporated by encoding the first frame with the CogVideoX encoder, adding random Gaussian noise (sampled independently from the video noise levels), projecting via a separate learnable ViT-like patchification layer, repeating sequence-wise to match video length, and summing with the video tokens. Training incorporates loss normalization and is conducted jointly on images and videos with variable resolutions (256, 512, and 1024), aspect ratios (16:9 and 9:16 for videos; 16:9, 9:16, and 1:1 for images), and video lengths (ranging from 17 to 385 frames). Videos are generated at 24 frames per second, and variable-FPS training is avoided due to observed performance decreases for target framerates without fine-tuning.
Base video DiT inference details.
Inference uses standard rectified flow without stochasticity. We find forty steps to balance quality and sampling speed effectively. For higher resolutions and longer video generation, a time-shifting strategy similar to Lumina-T2X is used, with a time shift of $\sqrt{32}$ for 1024-resolution videos.
Appendix C Training Details
[…]
Optimization details. We experiment with two model variants: a smaller generator with approximately 700 million parameters for ablations and initial explorations, and a larger 4 billion parameter model, which we use for the main results in this paper. Both models were trained with a batch size of 256 over 50,000 optimization steps with the LAMB optimizer (You et al., 2019). The learning rate was warmed up for the first 10,000 iterations from 0 to 0.005 and then linearly decreased to 0.0015 over subsequent iterations. For the large model, VD3D contains 230M trainable parameters in total which corresponds to around 5% of the total amount. Since the original video diffusion model is trained in the any-frame-conditioning pipeline (Menapace et al., 2024), we can produce variable camera trajectories from the same starting frame. For text conditioning, we use the T5-11B (Raffel et al., 2020) language model to encode text $1024$-dimensional embeddings into 128-length sequences. For training efficiency, they were precomputed for the entire dataset. The rest of the training details have been adopted from (Menapace et al., 2024).
Table 3: Multi-view generation. We evaluate all models using reference camera trajectories and single-view input images of the RealEstate10K test set. We compute reconstruction metrics based on the subsequent frames for the low-resolution and upsampled high-resolution generations.
Method Low-resolution High-resolution
PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Base Model 14.74 0.320 0.334 13.23 0.459 0.572
MotionCtrl (U-Net) 15.35 0.387 0.294 13.64 0.473 0.548
CameraCtrl (U-Net) 15.86 0.412 0.266 13.81 0.479 0.540
MotionCtrl 15.07 0.348 0.308 13.42 0.467 0.560
CameraCtrl 14.81 0.327 0.330 13.21 0.456 0.571
Ours 17.23 0.534 0.211 14.90 0.499 0.499
w/o Plucker 14.89 0.346 0.308 13.05 0.455 0.573
w/o ControlNet 14.66 0.313 0.340 13.10 0.450 0.573
w/o weight copy 16.96 0.509 0.220 14.75 0.495 0.504
w/o add context 14.45 0.303 0.368 13.06 0.446 0.588
w/o Plucker context 14.76 0.322 0.318 13.28 0.463 0.579
Table 4: Quality metrics evaluation. We evaluate all models using text prompts from the RealEstate10K and MSR-VTT test sets respectively.
Method RealEstate10K MSR-VTT
FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$) FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$)
Base Model 8.22 160.37 0.2677 3.50 141.26 0.2774
MotionCtrl (U-Net) 2.99 61.70 0.2646 16.85 283.12 0.2411
CameraCtrl (U-Net) 2.48 55.64 0.2681 12.33 201.33 0.2505
MotionCtrl 1.50 52.30 0.2708 9.97 183.57 0.2677
CameraCtrl 2.28 66.31 0.2730 8.47 181.90 0.2690
Ours 1.40 42.43 0.2807 7.80 165.18 0.2689
w/o Plucker 1.17 43.65 0.2715 9.84 152.91 0.2660
w/o ControlNet 3.66 137.06 0.2766 8.34 185.79 0.2674
w/o weight copy 1.38 42.00 0.2710 10.09 218.43 0.2647
w/o add context 1.45 44.74 0.2735 9.56 173.43 0.2657
w/o Plucker context 1.54 43.88 0.2724 8.83 168.95 0.2615
Table 5: Camera pose evaluation with additional baselines. We evaluate additional variants of the baselines using reference camera trajectories from the RealEstate10K test set. We compute translation and rotation errors based on estimated camera poses from generations using ParticleSfM (Zhao et al., 2022).
Method RealEstate10K MSR-VTT
MotionCtrl (latents) 0.549 0.183 0.650 0.145
CameraCtrl (patches) 0.587 0.197 0.648 0.233
MotionCtrl (frozen) 0.607 0.205 0.678 0.122
Ours 0.409 0.043 0.504 0.050
Table 6: Multi-view generation. We evaluate additional variants of the baselines using reference camera trajectories and single-view input images of the RealEstate10K test set. We compute reconstruction metrics based on the subsequent frames for the low-resolution and upsampled high-resolution generations.
Method Low-resolution High-resolution
PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
MotionCtrl (latents) 14.67 0.323 0.331 13.23 0.467 0.562
CameraCtrl (patches) 14.42 0.304 0.366 13.04 0.450 0.577
MotionCtrl (frozen) 14.59 0.308 0.340 13.11 0.455 0.573
Ours 17.23 0.534 0.211 14.90 0.499 0.499
Table 7: Quality metrics evaluation with additional baselines. We evaluate additional variants of the baselines using text prompts from the RealEstate10K and MSR-VTT test sets respectively.
Method RealEstate10K MSR-VTT
FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$) FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$)
MotionCtrl (latents) 1.83 77.39 0.2788 10.21 187.42 0.2636
CameraCtrl (patches) 2.57 71.04 0.2703 9.84 184.22 0.2612
MotionCtrl (frozen) 3.53 142.15 0.2772 8.19 165.48 0.2679
Ours 1.40 42.43 0.2807 7.80 165.18 0.2689
Table 8: Camera pose evaluation for random rotation trajectories. We evaluate all models using trajectories with randomly selected x-, y-, or z-axis and randomly sampled angular extent of the trajectory. We compute translation and rotation errors based on estimated camera poses from generations using ParticleSfM (Zhao et al., 2022).
Method RealEstate10K MSR-VTT
MotionCtrl 0.396 0.087 0.417 0.120
CameraCtrl 0.381 0.092 0.433 0.138
Ours 0.202 0.037 0.265 0.044
Table 9: Camera pose evaluation for random trajectories with translations and rotations. We evaluate all models using 11 trajectories for 1000 prompts with trajectories that involve significant directional changes including both rotations and translations. We compute translation and rotation errors based on estimated camera poses from generations using ParticleSfM (Zhao et al., 2022).
Method RealEstate10K MSR-VTT
MotionCtrl 0.451 0.095 0.456 0.146
CameraCtrl 0.369 0.088 0.479 0.135
Ours 0.236 0.041 0.258 0.050
Table 10: Camera pose evaluation with vanilla video DiT backbone. We incorporate our approach into a pre-trained vanilla DiT model in the latent space of CogVideoX (Yang et al., 2024d). We compute translation and rotation errors based on estimated camera poses from generations using ParticleSfM (Zhao et al., 2022).
Method RealEstate10K MSR-VTT
MotionCtrl 0.501 0.145 0.602 0.152
CameraCtrl 0.513 0.138 0.559 0.195
Ours 0.421 0.056 0.486 0.047
Table 11: Quality metrics evaluation with vanilla video DiT backbone. We incorporate our approach into a pre-trained vanilla DiT model in the latent space of CogVideoX (Yang et al., 2024d). We evaluate additional variants of the baselines using text prompts from the RealEstate10K and MSR-VTT test sets respectively.
Method RealEstate10K MSR-VTT
FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$) FID ($\downarrow$) FVD ($\downarrow$) CLIPSIM ($\uparrow$)
MotionCtrl 1.37 44.62 0.2752 9.68 157.90 0.2684
CameraCtrl 2.13 53.72 0.2748 8.32 152.88 0.2723
Ours 1.21 38.57 0.2834 6.88 137.62 0.2790